OpenStudy (anonymous):

Choosse the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1)

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

Well to find the slope use \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] but it requires you to find the point - slope form, which is \[ y-y _{1}=m(x-x _{1})\] so once you find the slope using the equation above you can plug it in as m on the point slope form equation, and your \[(x_1,y_1)\] points are the first coordinates you're given.

OpenStudy (anonymous):